# RmGPT: Rotating Machinery Generative Pretrained Model

### v1

@@ 4.4 Efficiency Analysis
4.5 Token Space Analysis
@@
5 Conclusion
This paper presents RmGPT, a foundation model for unified diagnosis and prognosis in rotating machinery. By utilizing a single model with consistent parameters across diverse datasets, RmGPT significantly outperforms SOTA algorithms, demonstrating exceptional accuracy and robustness in both diagnosis and prognosis tasks. The model ability to effectively handle varied equipment types, signal channels, fault categories and sampling frequencies underscores its adaptability and the feasibility of a unified approach in the PHM domain.
The result of few-shot learning experiments further highlight RmGPT capability to generalize from limited data, achieving remarkable performance even with minimal training samples. The token space analysis reveals that RmGPT transforms noisy input signals into a structured semantic space with clear class boundaries, indicating its proficiency in learning and distinguishing intrinsic fault patterns.
Future work will focus on extending RmGPT to a wider array of datasets, enhancing its generalization and scalability. Additionally, research into the local adaptation of large models will ensure their reliability and effectiveness in diverse operational contexts. These efforts aim to establish RmGPT as a universal diagnostic tool, advancing the field of PHM and paving the way for more adaptable and generalizable diagnostic models.
6 Acknowledgements

### v2

@@ IV-E Efficiency Analysis
IV-F Token Space Analysis
@@
V Conclusion
This paper presents RmGPT, a foundation model for unified diagnosis and prognosis in rotating machinery. By utilizing a single model with consistent parameters across diverse datasets, RmGPT significantly outperforms state-of-the-art (SOTA) algorithms, demonstrating exceptional accuracy and robustness in both diagnosis and prognosis tasks. The model’s ability to effectively handle varied equipment types, signal channels, fault categories, and sampling frequencies underscores its adaptability and validates the feasibility of a unified approach in the PHM domain.
The results of few-shot learning experiments further highlight RmGPT’s capability to generalize from limited data, achieving remarkable performance even with minimal training samples. The token space analysis reveals that RmGPT can transform noisy input signals into a structured semantic space with clear class boundaries, indicating its proficiency in learning and distinguishing intrinsic fault patterns.
This work represents a pioneering attempt to explore the feasibility of foundation models in the field of rotating machinery. RmGPT provides preliminary evidence that such models, which have shown great success in natural language and vision domains, can also be developed for complex industrial time-series scenarios. Our study not only validates the conceptual soundness of building a unified PHM foundation model, but also offers a practical framework that achieves superior performance across varied fault types and datasets.
Looking ahead, several important challenges and directions remain to be explored. From the perspective of depth, future research may investigate the relationship between model scale and downstream performance within the rotating machinery domain, and evaluate whether large-scale foundation models like RmGPT can effectively address the highly challenging problem of cross-device generalization. From the perspective of breadth, it would be promising to extend the generative token-based paradigm to other industrial time-series tasks—such as energy systems, batteries, wind turbines, or robotic health monitoring—thereby inspiring broader industrial applications of PHM foundation models.
Furthermore, we envision the development of a standardized capability evaluation framework for industrial foundation models, which would allow researchers to more systematically assess the robustness, adaptability, and scalability of such models under diverse industrial conditions. This would provide essential infrastructure for the next generation of intelligent and generalizable diagnostic systems in industry.
VI Acknowledgements